OpenStudy (anonymous):

Rhonda deposited $4,227.29 into a savings account with an interest rate of 4.9% compounded monthly. About how long will it take for the account to be worth $9,000?

OpenStudy (anonymous):

the interest rate you are given is annual but the interest is compounded monthly, what is the monthly interest rate?

OpenStudy (anonymous):

doest say thats al the info im givin

OpenStudy (anonymous):

no it doesn't say, but you can determine the monthly interest rate because you know how many months are in a year

OpenStudy (anonymous):

12

OpenStudy (anonymous):

yes. so how is that interest rate divided evenly amound the 12 months

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

\[4,227.29\times \left(1+\frac{.049}{12}\right)^{12t}=9000\] solve for \(t\)

OpenStudy (anonymous):

what do i do with the 12t?